OpenStudy (anonymous):

the volume of rectangle prism can computer using the formula v=lwh. find the length of a prism that has a volume of 1344 cubic cm , a width of 8 cm , and a height of 12 cm

OpenStudy (campbell_st):

just substitute the given information and then solve for l \[1344 = l \times 8 \times 12~~~~~or~~~~~~1344 = l \times 96\]